OpenStudy (anonymous):

You roll a blue number cube and a green number cube. Find P a number less than 2 on the blue cube and an odd number on the green cube.

OpenStudy (cwrw238):

these are independent events so you multiply the individual properties P(<2) = 1/6 P(odd number) = 1/2

OpenStudy (cwrw238):

so Probability required = 1/2 * 1/6

OpenStudy (cwrw238):

ok?

OpenStudy (anonymous):

Yea. I got it . I think.

OpenStudy (cwrw238):

1/12 right?

OpenStudy (anonymous):

Yep that's what i had.